OpenStudy (anonymous):

The expression 14^6 ÷ 14^4 is equal to _____. 14^2 2 14^10 1

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{a^6}{a^4}={a^6}{a^{-4}}=a^{6-4} =a^{2}\]

OpenStudy (fibonaccichick666):

Would you like to work through it @Susan_Hester ?

OpenStudy (anonymous):

yes please

OpenStudy (fibonaccichick666):

ok so let's discuss what an exponent represents first.

OpenStudy (fibonaccichick666):

So just like multiplication is shorthand for addition, exponents are shorthand for multiplication

OpenStudy (fibonaccichick666):

does that make sense?

OpenStudy (anonymous):

yeah

OpenStudy (fibonaccichick666):

ok, so if we have \(2^3=2*2*2\) do you follow?

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

ok so now, let's do a similar problem to yours: \[3^9 \div 3^6\] Now first thing I am going to do here is write this as a fraction, I like them better. \[\frac{3^9}{3^6}\]

OpenStudy (fibonaccichick666):

so, if we wanted to do this the long but make sense way, we will expand 3^9 and 3^6 like so \[\frac{3*3*3*3*3*3*3*3*3}{3*3*3*3*3*3}\]

OpenStudy (fibonaccichick666):

Do you understand to here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

? u there

OpenStudy (fibonaccichick666):

so now, can you simplify the fraction?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

9/6

OpenStudy (fibonaccichick666):

it gets even simpler

OpenStudy (anonymous):

3/2

OpenStudy (fibonaccichick666):

how did you get 6 on the bottom?

OpenStudy (anonymous):

i counted 6

OpenStudy (fibonaccichick666):

ahhh, ok, no you can't do it that way, think about if you had \[\frac{12}{6}\] how would you simplify that?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

12 divided by 6 is 2?

OpenStudy (fibonaccichick666):

yea, so we are doing the same thing with all that multiplication

OpenStudy (fibonaccichick666):

I could write 12/6 as \[\frac{3*2*2}{2*3}\]

OpenStudy (anonymous):

need it in exponet form

OpenStudy (fibonaccichick666):

I need to make 1/1 so because multiplication is commutative, I can pull the 3s, and 2s. (ie cancel them) \[\frac{\not3*\not 2*2}{\not 3*\not2}\] So I am left with only 2

OpenStudy (fibonaccichick666):

now back to our exponent example, we use the same process, try one more time please \[\frac{3*3*3*3*3*3*3*3*3}{3*3*3*3*3*3}\]

OpenStudy (anonymous):

14*14*\[14*14*14*14*14*14/14*14*14*14\]

OpenStudy (anonymous):

can u just tell me the answer to my question please

OpenStudy (anonymous):

ill give u a medal

OpenStudy (anonymous):

?

OpenStudy (fibonaccichick666):

You can do it just like the one above, or you can memorize the exponent rules found here. http://www.purplemath.com/modules/exponent.htm We were simply going to derive them so they make sense

OpenStudy (anonymous):

never mind!